OpenStudy (anonymous):

A card is drawn from a standard deck of cards. Determine whether the events are mutually exclusive or inclusive. Then find the probability. 1.P(queen or spade)

OpenStudy (anonymous):

there total of 13 spades cards and 4 queen but one card is common in spade and queen so, P(queen or spade)= 13/52+4/52-1/52= 16/52= 8/27

OpenStudy (anonymous):

good job,ty